(1) A sales tax of $6.18 is added to a purchase of $95. Find the tax rate.

(2) Bob receives 3.5% commission on his sales job. Find the commission if his sales total $685.

(3) Louis earned $2375 on a sale of $95,000. Find his commission rate.

(4) A computer is on sale for 15% off the regular price of $2800. Find the discount and the sale price.

Answer by checkley71(8403)

6.18/95.00=6.5%
685*.035=23.975
2375/95000=2.5%
2800*.15=420 WHICH IS THE DISCOUNT THUS THE SALE PRICE IS 2800-420=2380